# Reinforced Generation of Combinatorial Structures: Applications to Complexity Theory

### v4

@@ 2 AlphaEvolve as a Framework for Combinatorial Discovery
AlphaEvolve [30, 34] is a code-mutation LLM agent capable of autonomously evolving a code snippet to hill-climb on a space of objects with a well defined evaluation metric. We use AlphaEvolve, with problem-specific verifiers that score snippets generated by AlphaEvolve, to discover novel finite combinatorial structures that improve results in complexity theory [24, 14, 26]. Previously AlphaEvolve has been used in a variety of scientific domains [34, 30]. In the following, we provide a self-contained description of AlphaEvolve and the verifier. Additionally, we highlight the system enhancements needed to achieve the combinatorial structures in this paper.
@@
2.1 Propose-test-refine (PTR) paradigm
@@
    Suitably defined conjecture/claim: The first step towards discovery is a concrete definition of the combinatorial structure being sought. For example, it can be a weighted graph with some fixed set of nodes, or a SAT formula with a fixed number of clauses and variables. In this paper, we restrict these to finite structures rather than a parameterized family  [^4] Let $\mathcal{S}$ be a class of combinatorial structures (e.g., graphs, hypergraphs, codes, formulas). A parameterized family of structures from $\mathcal{S}$ is a sequence $\{S_{n}\}_{n\in\mathbb{N}}$ s.t. each $S_{n}\in\mathcal{S}$ is a finite structure whose size depends on the parameter $n$.. While the source and the target problems form a parameterized family (which enables us to prove theorems with $\forall n$ quantification), the gadgets used to translate between them are finite structures. (A gadget may be thought of as a combinatorial structure mapping a source constraint to a linear combination of target constraints [19, 36].)
@@
    An initial candidate search algorithm: AlphaEvolve operates by making modifications to a code-snippet by asking an LLM to make syntactically valid modifications, while improving some fitness score. Denote the code snippet at time $t$ by $\mathbf{C}_{t}:\mathcal{D}\to\mathcal{S}$, where $\mathcal{D}$ is the domain capturing problem specific parameters and any inputs to the snippet, and $\mathcal{S}$ is the range space of all possible combinatorial structures. Additionally, let $\texttt{score}:\mathcal{S}\to\mathbb{R}$ be the function that assigns a fitness score to a structure. The next function $\mathbf{C}_{t+1}$ is chosen as follows:
    $\mathbf{C}_{t+1}\leftarrow\texttt{AlphaEvolve}\left(\left(\mathbf{C}_{0}(d_{0}),\texttt{score}(\mathbf{C}_{0}(d_{0}))\right),\ldots,\left(\mathbf{C}_{t}(d_{t}),\texttt{score}(\mathbf{C}_{t}(d_{t}))\right)\right)$ (1)
    AlphaEvolve makes an LLM call using the tuple of inputs in the prompt to the LLM, prompting it to output the description of $\mathbf{C}_{t+1}$, with the goal of increasing the score [^5] Technically, AlphaEvolve does not feed in all the prior functions as the prompt to the LLM. Instead, it uses a genetic algorithm to select a subset of them while maintaining high scores, and diversity [34, Figure 1].. We emphasize the following subtle point here: AlphaEvolve does not directly search for combinatorial structures, but instead  a code-snippet that would output a combinatorial structure with a high score. Figure 2 shows an initial code-snippet to be evolved by AlphaEvolve.
@@
    A verifier with a smooth scoring function: In the problems we consider, we seek combinatorial structures in spaces of size exponential in the description of the structure. As a result, a well-defined verifier that validates (and scores) the constructions generated by the $\mathbf{C}_{t}$’s (in Equation 1) is necessary. In particular, the verifier serves as a constraint in the optimization landscape.
@@
    Concretely, for average-case hardness in Section 3, this corresponds to validating whether the graphs are Ramanujan with certain properties (e.g., having a large cut, or independent set), and for worst-case NP-hardness in Section 4, this corresponds to verifying whether the soundness and completeness constraints (e.g., [36, 19]) corresponding to the gadget reduction are satisfied.
@@
    As mentioned earlier, every structure generated by AlphaEvolve needs to be scored, including invalid ones that violate certain constraints. The scoring function $\texttt{score}:\mathcal{S}\to\mathbb{R}$ is used to hill-climb on the optimization landscape. For invalid structures, it decides how far from the constraint boundary the current structure lies. (A scoring function is smooth if its value gracefully increases/decreases as one moves away from the constraint boundary.) This is where the most creativity, and problem specific domain expertise comes in. In particular, the efficacy of the scoring function largely decides the overall capability of the system to discover novel structures. Also, scoring is usually the most compute intensive step in the PTR paradigm.

### v5

@@ Appendix A AlphaEvolve as a Framework for Combinatorial Discovery
AlphaEvolve [42, 48] is an LLM-based code-mutation agent. In the context of this paper, it evolves code snippets that generate combinatorial structures, scoring generated structures by their fitness for the problem at at hand. These evolved code snippets generated novel finite combinatorial structures that improve results in the hardness of approximation. Previously AlphaEvolve has been used in other of scientific domains [48, 42]. In the following, we provide a self-contained description of AlphaEvolve. Additionally, we highlight the system enhancements needed to achieve the combinatorial structures in this paper.
@@
A.1 Propose-test-refine (PTR) paradigm
@@
    Suitably defined conjecture/claim: The first step towards discovery is a concrete definition of the combinatorial structure being sought. For example, it can be a weighted graph with some fixed set of vertices, or a SAT formula with a fixed number of clauses and variables. In this paper, we restrict these to finite structures rather than a parameterized family  [^11] Let $\mathcal{S}$ be a class of combinatorial structures (e.g., graphs, hypergraphs, codes, formulas). A parameterized family of structures from $\mathcal{S}$ is a sequence $\{S_{n}\}_{n\in\mathbb{N}}$ s.t. each $S_{n}\in\mathcal{S}$ is a finite structure whose size depends on the parameter $n$.. While the source and the target problems form a parameterized family (which enables us to prove theorems with $\forall n$ quantification), the gadgets used to translate between them are finite structures. (A gadget may be thought of as a combinatorial structure mapping a source constraint to a linear combination of target constraints [26, 50].)
@@
    An initial candidate search algorithm: AlphaEvolve uses an LLM to make syntactically valid modifications to a code snippet, with the goal of improving a fitness score for the combinatorial structure it generates. Denote the code snippet at time $t$ by $\mathcal{C}_{t}:\mathcal{D}\to\mathcal{S}$, where $\mathcal{D}$ is the domain capturing problem specific parameters and any inputs to the snippet, and $\mathcal{S}$ is the range space of all possible combinatorial structures. Additionally, let $\texttt{score}:\mathcal{S}\to\mathbb{R}$ be the function that assigns a fitness score to a structure. The next function $\mathcal{C}_{t+1}$ is chosen as follows:
    $\mathcal{C}_{t+1}\leftarrow\texttt{AlphaEvolve}\left(\left(\mathcal{C}_{0}(d_{0}),\texttt{score}(\mathcal{C}_{0}(d_{0}))\right),\ldots,\left(\mathcal{C}_{t}(d_{t}),\texttt{score}(\mathcal{C}_{t}(d_{t}))\right)\right)$ (1)
    AlphaEvolve makes an LLM call using the tuple of inputs in the prompt to the LLM, prompting it to output the description of $\mathcal{C}_{t+1}$, with the goal of increasing the score [^12] Technically, AlphaEvolve does not feed in all the prior functions as the prompt to the LLM. Instead, it uses a genetic algorithm to select a subset of them while maintaining high scores, and diversity [48, Figure 1].. We emphasize the following subtle point: AlphaEvolve does not directly search for combinatorial structures, but instead for a code snippet that would output a combinatorial structure with a high score. Figure 6 shows an initial code-snippet to be evolved by AlphaEvolve.
@@
    A verifier with a smooth scoring function: In the problems we consider, we seek combinatorial structures in spaces of size exponential in the description of the structure. A well-defined verifier that validates (and scores) the constructions generated by the $\mathcal{C}_{t}$’s (in Equation 1) guides this search.
@@
    Concretely, for average-case hardness in Section 2, this corresponds to validating whether the graphs are Ramanujan with certain properties (e.g., having a large cut, or independent set), and for worst-case NP-hardness in Sections 3 and 4, this corresponds to verifying whether the soundness and completeness constraints (e.g., [50, 26]Lemmas D.3 and D.2) for the gadget reduction are satisfied.
@@
    As mentioned earlier, every structure generated by AlphaEvolve needs to be scored, including invalid ones that violate certain constraints. The scoring function $\texttt{score}:\mathcal{S}\to\mathbb{R}$ is used to hill-climb on a optimization landscape over the combinatorial structures. For invalid structures, it decides how far from the constraint boundary the current structure lies. (Informally, a scoring function is smooth if its value gracefully increases/decreases as one moves away from the constraint boundary.) This is where the most creativity and problem-specific domain expertise comes in. In particular, the efficacy of the scoring function largely decides the overall capability of the system to discover novel structures. Also, scoring is usually the most compute intensive step in the PTR paradigm.